# Spatial and temporal cluster tomography of active matter

### v1

@@ Abstract
Figure 1: Cluster tomography. (a) Spatial cluster tomography. Given a cluster-forming system, we consider one-dimensional cross-sections and tally the distance $s$ between successive observations of each cluster, to detect and characterize phase transitions. (b) Temporal cluster tomography. Given a pair of particles (red), we measure the time intervals $s_{t}$ between successive observations of them belonging to the same cluster (first and third snapshots) to study the dynamics of the system.
@@
As discussed in detail in App. A, a non-zero lacunarity leads to the special case of a logarithmic “corner contribution” $\mathcal{C}=b\ln\ell$ in the number of clusters intersecting a line of length $\ell$ at criticality [19, 32]. The corner contribution has been studied in a number of equilibrium classical [29, 30, 31] and quantum [20, 33, 34, 12] systems in the vicinity of a critical point for the more general case of contour lines with various corner angles—e.g. squares and triangles (a line segment corresponds to two corners with $2\pi$ angles). Here, by means of the gap-size statistics, we demonstrate that corner contributions also yield meaningful insight in active systems.
@@
Concretely, we consider two-dimensional ABPs, i.e. $N$ polar particles with positions $\mathbf{r}_{i}$ and polarities $\hat{\mathbf{n}}_{i}=(\cos\theta_{i},\sin\theta_{i})$ whose dynamics follow
$\partial_{t}{\mathbf{r}_{i}}$ $=v_{0}\hat{\mathbf{n}}_{i}(\theta_{i})+\mu\sum_{j\neq i}\mathbf{F}_{ij}(r_{ij})\;,$ $\partial_{t}{\theta_{i}}$ $=\sqrt{2D_{r}}\,\xi_{i}(t)\;,$ (3)
where $v_{0}$ is the self-propulsion speed, $\mu$ is the mobility coefficient, $D_{r}$ is the rotational diffusion coefficient, and $\xi_{i}$ denotes zero-mean, unit-variance Gaussian noise. We set inter-particle forces to be finite-range harmonic repulsion, $\mathbf{F}_{ij}=k(a-r_{ij})\hat{\mathbf{r}}_{ij}$ if $r_{ij}<a$, and 0 otherwise, where $k$ is a stiffness constant, $a$ is the diameter of the repulsive disks, $|r_{ij}=\absolutevalue{\mathbf{r}_{i}-\mathbf{r}_{j}}$, and $\hat{\mathbf{r}}_{ij}=(\mathbf{r}_{i}-\mathbf{r}_{j})/r_{ij}$. We henceforth set the softness of the repulsion via $\mu ka/v_{0}=10^{2}$. We simulate Eq. 3 in a square box with periodic boundary conditions and side-length $L$. Two dimensionless parameters determine together whether MIPS is observed [46, 47]: the volume fraction occupied by the disks, $\phi=N\pi a^{2}/L^{2}$, and the ratio of the self-propulsion and rotational diffusion time-scales, often called rotational Péclet number, $\mathrm{Pe}_{r}=v_{0}/(aD_{r})$. In the $(\phi,\mathrm{Pe}_{r})$ plane, MIPS is observed in a U-shaped region at large angular Péclet numbers and intermediate volume fractions [47] (see SM [[SeeSupplementalMaterialat][foracompletedescriptionofnumericalmethods, afullphasediagramofourchoiceofABPdynamics, andadditionalmethodologicalnotesonclustertomography.]supp]).
@@
Figure 2: Spatial cluster tomography. (a) Gap-size statistics at $\mathrm{Pe}_{r}=10$ and volume fractions below, close to, and above the critical region. (b) Fraction of the available volume occupied by the largest cluster. (c) Corner contribution $\mathcal{C}$, excluding the largest cluster. The dashed line marks $\phi^{*}=0.515$. (Inset) Data collapse of the corner contribution using the Ising values $b=0.0608$, $\nu=1$. $L\leq 32$ data are omitted for clarity. (d) Finite-size estimates of the lacunarity using $\ell=L/8$. An infinite-size extrapolation in $\ln\ell/\ell$ shows consistency with the Ising lacunarity (dashed horizontal line). Where no error bars are visible, the uncertainty is smaller than the marker.
@@
We now study the lacunarity $b$ to determine the universality class of the critical point. We extract finite-size estimates of $b$ from the gap-size statistics using (App. A)
@@
Figure 3: Temporal cluster tomography. (a) Representative time series for a particle pair at volume fractions below, at, and above the critical region over 200 time steps ($L=512$). Vertical lines mark snapshots where the pair is in the same cluster. (b) Temporal gap-size statistics at the same volume fractions as (a), normalized by the number of snapshots $T$. (c) Burstiness parameter computed from $g_{t}(s_{t})$. The dashed line marks $\phi^{*}=0.515$. Where no error bars are visible, the uncertainty is smaller than the marker.
@@
The temporal gap-size statistics $g_{t}(s_{t})$, shown in Fig. 3, displays a qualitative change close to $\phi^{*}=0.515$. At low volume fractions, $g_{t}(s_{t})$ is found to be heavy-tailed, indicating that large temporal gaps are common, especially as the system size increases. Above $\phi^{*}$, only short gaps are present, and increasing the system size does not alter the observed statistics (Fig. 3). This qualitative change in the dynamical behavior translates into a rapid drop in the burstiness parameter $B$, from a positive to a negative value across $\phi^{*}$ (Fig. 3).
@@
Our results can be interpreted in light of the correlation length of the system. Below $\phi^{*}$, cluster size is bounded by the finite correlation length $\xi$. As long as the distance between two particles is larger than $\xi$, they are unlikely to be found in the same cluster. However, if the distance between them becomes comparable to $\xi$, they will often switch between belonging to the same cluster then to different ones, as clusters continuously form, merge, and break apart. We observe this as sequences of short temporal gaps separated by long temporal gaps, typical of bursty ($B>0$) dynamics (Fig. 3). On the other hand, above $\phi^{*}$, the system essentially contains a single cluster with empty pockets. When a particle leaves the cluster, the typical gap time is set by the typical size of empty regions in the system and the self-propulsion speed. As a result, the frequency of temporal gaps drops sharply after $s_{t}=1$, leading to $B<0$.
In the critical region, since the typical time-scale of order parameter fluctuations diverges [49], one could naïvely expect long temporal gaps to be common, resulting in a large positive $B$. Surprisingly, we instead find that the burstiness parameter is negative. More surprising yet, its magnitude decreases with system size: as $L\to\infty$, the burstiness parameter approaches $B=0$, a value typical of uncorrelated, memory-less dynamics [24]—in sharp contrast to the highly-correlated dynamics of the order parameter. In other words, the dynamics of cluster membership are largely independent from order parameter fluctuations, and seem to be governed by different physics than the standard theory of dynamic critical phenomena [49].
@@
First-order transition—Finally, we demonstrate that cluster tomography is not limited to the study of continuous phase-transitions. To do so, we carry out a similar analysis on simulations with $\mathrm{Pe}_{r}=100$, much larger than the value at the MIPS critical point at $\mathrm{Pe}_{r}\approx 10$ [47, 48]. As we vary the volume fraction between $\phi=0.2$ and $0.3$, we see a sudden qualitative change in the shape of the gap-size statistics and a corresponding jump in the corner contribution (Fig. 4 and 4), consistent with the presence of a first-order phase transition  [^1] Note that the largest cluster was included in this analysis, as its exclusion was only helpful to mitigating finite-size effects close to the critical point.. We observe a similar sudden change in the shape of the temporal gap-size statistics, and a corresponding jump in the burstiness parameter, marking the transition to a different dynamical regime (Fig. 4 and 4).
The physics at high-$\mathrm{Pe}_{r}$ results in a connection between spatial and temporal gap-size statistics—in the phase-separated regime ($\phi\geq 0.25$), they seem to have roughly the same shape (Fig. 4 and 4). To understand this, notice that at high $\mathrm{Pe}_{r}$, particles leaving a cluster travel in a nearly straight line before hitting another cluster. Since the system is dominated by one large cluster, $g(s)$ roughly gives the distribution of traveled distances. Altogether, the duration of a temporal gap is well approximated by $s/v_{0}$, and $g_{t}(s_{t})$ is approximately proportional to $\hat{g}(s)$. This proportionality is remarkable, as spatial and temporal cluster tomography generally provide independent information. In binary random media, such a connection between geometry and dynamics is only observed in the limit of Knudsen diffusion [26].
@@
Figure 4: First-order transition. Gap-size statistics (a), corner contribution (b), temporal gap-size statistics (c), and burstiness parameter (d) across the MIPS boundary at $\mathrm{Pe}_{r}=100$, $L=256$.
@@
We also introduced temporal cluster tomography to probe dynamical properties across phases. By analogy with complex networks, we introduced the burstiness parameter $B$ [24, 25], an informative and easily accessible summary statistics that can be used to characterize active dynamics. We reported a heavy-tailed, bursty ($B>0$) distribution of temporal gaps in the sparse (low–volume fraction) phase, and a more regular ($B<0$) dynamics in the high-volume fraction phase—with, surprisingly, $B\approx 0$ at the critical point.
@@ Acknowledgements.
This work was supported by the National Science Foundation under Grant No. PHY-2310706 of the QIS program in the Division of Physics, and by the Baker Faculty Grant of the Weinberg College of Arts and Sciences, Northwestern University. We thank Monica Olvera de la Cruz, Daniel Matoz-Fernandez, Michelle Driscoll, Brendan Blackwell, Helen S. Ansell, Sean P. E. Dougherty, and Xiangyi Meng for useful discussions.

### v2

@@ Abstract
FIG. 1: Cluster tomography. (a) Spatial cluster tomography. Given a cluster-forming system, we consider one-dimensional cross-sections and tally the distance $s$ between successive observations of each cluster, to detect and characterize phase transitions. (b) Temporal cluster tomography. Given a pair of particles (red), we measure the time intervals $s_{t}$ between successive observations of them belonging to the same cluster (first and third snapshots) to study the dynamics of the system.
@@
As discussed in detail in Appendix A, a non-zero lacunarity leads to the special case of a logarithmic “corner contribution” $\mathcal{C}=b\ln\ell$ in the number of clusters intersecting a line of length $\ell$ at criticality [19, 32]. The corner contribution has been studied in a number of equilibrium classical [29, 30, 31] and quantum [20, 33, 34, 12] systems in the vicinity of a critical point for the more general case of contour lines with various corner angles—e.g. squares and triangles (a line segment corresponds to two corners with $2\pi$ angles). Here, by means of the gap-size statistics, we demonstrate that corner contributions also yield meaningful insight in active systems.
@@
Concretely, we consider two-dimensional ABPs, i.e. polar particles with diameter $a$ and self-propulsion speed $v_{0}$, interacting through finite-range harmonic repulsion and subject to rotational noise. A full description of our choice of ABP dynamics can be found in Appendix B. We simulate $N$ particles in a square box with periodic boundary conditions and side-length $L$. Two dimensionless parameters determine together whether MIPS is observed [46, 47]: the volume fraction occupied by the disks, $\phi=N\pi a^{2}/L^{2}$, and the ratio of the self-propulsion and rotational diffusion time-scales, often called rotational Péclet number, $\mathrm{Pe}_{r}=v_{0}/(aD_{r})$, where $D_{r}$ is the rotational diffusion coefficient. In the $(\phi,\mathrm{Pe}_{r})$ plane, MIPS is observed in a U-shaped region at large angular Péclet numbers and intermediate volume fractions [47]. See Supplementary Material (SM) for a full phase diagram [[SeeSupplementalMaterialat][foracompletedescriptionofnumericalmethods, afullphasediagramofourchoiceofABPdynamics, andadditionalnotesontheimplementationofclustertomography.]supp].
@@
FIG. 2: Spatial cluster tomography. (a) Gap-size statistics at $\mathrm{Pe}_{r}=10$ and volume fractions below, close to, and above the critical region. (b) Fraction of the available volume occupied by the largest cluster. (c) Corner contribution $\mathcal{C}$, excluding the largest cluster. The dashed line marks $\phi^{*}=0.515$. (Inset) Data collapse of the corner contribution using the Ising values $b=0.0608$, $\nu=1$. $L\leq 32$ data are omitted for clarity. (d) Finite-size estimates of the lacunarity using $\ell=L/8$. An infinite-size extrapolation in $\ln\ell/\ell$ shows consistency with the Ising lacunarity (dashed horizontal line). Where no error bars are visible, the uncertainty is smaller than the marker.
@@
We now study the lacunarity $b$ to determine the universality class of the critical point. We extract finite-size estimates of $b$ from the gap-size statistics using (Appendix A)
@@
FIG. 3: Temporal cluster tomography. (a) Representative time series for a particle pair at volume fractions below, at, and above the critical region over 200 time steps ($L=512$). Vertical lines mark snapshots where the pair is in the same cluster. (b) Temporal gap-size statistics at the same volume fractions as (a), normalized by the number of snapshots $T$. (c) Burstiness parameter calculated from $g_{t}(s_{t})$. To avoid artefacts due to the periodic boundaries, only temporal gaps $s_{t}<L$ are used in the calculation. The dashed line marks $\phi^{*}=0.515$. Where no error bars are visible, the uncertainty is smaller than the marker.
@@
The temporal gap-size statistics $g_{t}(s_{t})$, shown in Fig. 3, displays a clear qualitative change close to $\phi^{*}=0.515$. At low volume fractions, $g_{t}(s_{t})$ is found to be heavy-tailed, indicating that large temporal gaps are common, especially as the system size increases. Above $\phi^{*}$, only short gaps are present, and increasing the system size does not alter the observed statistics (Fig. 3). This qualitative change in the dynamics translates into a shift in the burstiness parameter $B$ from a positive to a negative value near $\phi^{*}$ (Fig. 3). At low volume fractions, we find a large positive $B$ that increases with the system size and remains approximately constant for varying $\phi$. In contrast, above $\phi^{*}$, $B$ exhibits a strong dependence on $\phi$ but not on $L$ and quickly becomes negative at higher volume fractions.
@@
Our results can be interpreted in light of the correlation length of the system. Below $\phi^{*}$, cluster size is bounded by the finite correlation length $\xi$. As long as the distance between two particles is larger than $\xi$, they are unlikely to be found in the same cluster. However, if the distance between them becomes comparable to $\xi$, they will often switch between belonging to the same cluster then to different ones, as clusters continuously form, merge, and break apart. We observe this as sequences of short temporal gaps separated by long temporal gaps, typical of bursty ($B>0$) dynamics (Fig. 3). On the other hand, when $\phi\gg\phi^{*}$, the system essentially contains a single cluster with empty pockets. When a particle leaves the cluster, the typical gap time is set by the typical size of empty regions in the system and the self-propulsion speed. As a result, the frequency of temporal gaps drops sharply after $s_{t}=1$, leading to $B<0$.
@@
First-order transition—Finally, we demonstrate that cluster tomography is not limited to the study of continuous phase-transitions. To do so, we carry out a similar analysis on simulations with $\mathrm{Pe}_{r}=100$, much larger than the value at the MIPS critical point at $\mathrm{Pe}_{r}\approx 10$ [47, 48]. As we vary the volume fraction between $\phi=0.2$ and $0.3$, we see a sudden qualitative change in the shape of the gap-size statistics and a corresponding jump in the corner contribution (Figs. 4 and 4), consistent with the presence of a first-order phase transition  [^1] Note that the largest cluster was included in this analysis, as its exclusion was only helpful to mitigating finite-size effects close to the critical point.. We observe a similar sudden change in the shape of the temporal gap-size statistics, and a corresponding jump in the burstiness parameter, marking the transition to a different dynamical regime (Figs. 4 and 4). We note that, in the phase-separated regime, $B$ is sensitive to changes in the cluster structure—e.g. the largest cluster temporarily splitting in two—leading to a larger variability, not entirely accounted for by our estimate of the uncertainty.
@@
FIG. 4: First-order transition. Gap-size statistics (a), corner contribution (b), temporal gap-size statistics (c), and burstiness parameter (d) across the MIPS boundary at $\mathrm{Pe}_{r}=100$, $L=256$.
@@
We also introduced temporal cluster tomography to probe dynamical properties across phases. By analogy with complex networks, we introduced the burstiness parameter $B$ [24, 25], an informative and easily accessible summary statistics that can be used to characterize active dynamics. We reported a heavy-tailed, bursty ($B>0$) distribution of temporal gaps in the sparse (low–volume fraction) phase, and a more regular ($B<0$) dynamics in the high-volume fraction phase. While we observed $B\approx 0$ at intermediate volume fractions, this value does not appear to correspond to a physically significant point in parameter space, which highlights a potential limitation of the burstiness parameter as a summary statistics for temporal cluster tomography. Future work could identify alternative metrics to capture the rich information contained in the temporal gap-size statistics.
@@ Acknowledgements.
This work was supported by the National Science Foundation under Grant No. PHY-2310706 of the QIS program in the Division of Physics, and by the Baker Faculty Grant of the Weinberg College of Arts and Sciences, Northwestern University. This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University, which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. We thank Monica Olvera de la Cruz, Daniel Matoz-Fernandez, Michelle Driscoll, Brendan Blackwell, Helen S. Ansell, Sean P. E. Dougherty, and Xiangyi Meng for useful discussions.
@@ Appendix A Gap sizes and the corner contribution
Appendix B Implementation of ABP dynamics
Our model of ABPs consists of polar particles with positions $\mathbf{r}_{i}$ and polarities $\hat{\mathbf{n}}_{i}=(\cos\theta_{i},\sin\theta_{i})$ whose dynamics follow
$\partial_{t}{\mathbf{r}_{i}}$ $=v_{0}\hat{\mathbf{n}}_{i}(\theta_{i})+\mu\sum_{j\neq i}\mathbf{F}_{ij}(r_{ij})\;,$ $\partial_{t}{\theta_{i}}$ $=\sqrt{2D_{r}}\,\xi_{i}(t)\;,$ (B1)
where $v_{0}$ is the self-propulsion speed, $\mu$ is the mobility coefficient, $D_{r}$ is the rotational diffusion coefficient, and $\xi_{i}$ denotes zero-mean, unit-variance Gaussian noise. We set inter-particle forces to be finite-range harmonic repulsion, $\mathbf{F}_{ij}=k(a-r_{ij})\hat{\mathbf{r}}_{ij}$ if $r_{ij}<a$, and 0 otherwise, where $k$ is a stiffness constant, $a$ is the diameter of the repulsive disks, $|r_{ij}=\absolutevalue{\mathbf{r}_{i}-\mathbf{r}_{j}}$, and $\hat{\mathbf{r}}_{ij}=(\mathbf{r}_{i}-\mathbf{r}_{j})/r_{ij}$. We set the softness of the repulsion via $\mu ka/v_{0}=10^{2}$. A detailed discussion of our numerical simulations of Eq. B1 can be found in the SM [48].